17. Blue light has a wavelength of 4.75 X 100 m. What is the frequency when it is moving at 3 X 109 m/s (2 pts)
Answer #1

Solution)

Given,

Wavelength, lamda=4.75*10^-9 m

Speed, v=3*10^8 m/s

We know,

Velocity, v= f*lambda

So, frequency, f=(3*10^8)/(4.75*10^-9)= 6.31*10^16 Hz (Ans)

=============